Myrna Cruz invested a certain amount at Family Bank at 14% interest and another amount in government bonds at 12%, with both investment giving her annual income of $2420. Had she invested both amounts in government bonds, she would have realized only $2100 annually. How much did she invest at each rate?
Answer by richwmiller(17219)

0.12*g+0.14*f=2420,
.12(g+f)=2100
0.12*g+0.14*f=2420,
0.12*g+0.12*f=2100
.02f=320
f=16000
g=1500